Mission Manufacturing’s depreciation expense for the most recent year was $72,000 and its net income was $424,000. The company’s net cash flow from operating activities was $569,600. Other information from Mission’s financial statements includes a decrease in accounts receivable of $24,000, a $27,200 increase in accounts payable, and a $9,600 decrease in prepaid expenses. The only other activity that occurred for the company this year was the sale of a piece of equipment whose book value was $16,000. How much was the gain or loss on this sale?
Answer #1
We have the equation
Cash flow from operating activites = Net income-Gain on sale of equipment+Depreciation+Decrease in accounts receivable+increase in accounts payable+Decrease in prepaid expenses
Substituting available values, we have:
569600 = 424000+72000+24000+27200+9600+Gain on sale of equipment
Therefore gain on sale of equipment = 569600 -(424000+72000+24000+27200+9600) = $    12,800